OpenStudy (anonymous):

-2(v-7)-6v=8(4-v)-3 step by step plz

OpenStudy (anonymous):

Same approach: -2*v - (-2)*7 - 6v = 8*4 - 8*v - 3 -2v + 14 - 6v = 32 - 8v - 3 BTW - are you sure the problem is right? I am not getting any value of v here.

OpenStudy (anonymous):

yeah its right

OpenStudy (anonymous):

Well, something is wrong as you get -8v on both sides that eliminates v.

OpenStudy (anonymous):

it sys it is right inmy book

OpenStudy (anonymous):

Don't know why. But, there is no solution to this equation.

OpenStudy (anonymous):

then the answer would be no solution

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

okay thank you